# Benchmarking Differentially Private Tabular Data Synthesis

### v1

@@ 9.3. Module Comparison
Figure 6. The maximum attribute domain size of Loan dataset under fixed merging thresholds and different $\varepsilon$. In the figure, “$0.0\%$” means no fixed threshold (merge only by $3\sigma$ criteria), and “None” means no merging preprocessing (raw data).
Figure 7. Average scaled metrics for different selection modules on different datasets. We use PGM as the common synthesizer and match PGM with different marginal selection modules. The machine learning efficiency, query error, and fidelity error are scaled using ground truth values provided in Table 2 while running time is scaled using a simple min-max linear normalization. The small dataset includes ACSincome and ACSemploy, whereas the large dataset consists of Bank, Higgs-small, and Loan.
Figure 8. Average scaled metrics for different synthesis modules on different datasets. We use PrivSyn as the common marginal selector and try different synthesizers. The machine learning efficiency, query error, and fidelity error are scaled using ground truth values provided in Table 2 while running time is scaled using a simple min-max linear normalization. The small datasets include ACSincome and ACSemploy, whereas the large datasets consist of Bank, Higgs-small, and Loan.
@@
Comparison of Synthesis Modules. We compare the synthesis algorithms of all six algorithms: GUM, PGM, relaxed projection (RP), genetic algorithm (GA), and generative network (GN). PGM is employed by both AIM and PrivMRF. To avoid multiple fitting steps from adaptive selection and control the running time, we use the PrivSyn selection algorithm as the common marginal selector. The results are shown in Figure 8. GUM and PGM are conducted on CPUs, and the remaining works are on GPUs.
Among these methods, PGM and the generative network achieve the best overall accuracy and stability in fitting features on small datasets. PGM benefits from its expressive structure, which, as indicated by Equation 2, infers marginals without refitting existing ones, thus minimizing compounding errors. However, the PGM’s fitting process is quite slow, due to the densely selected marginals by PrivSyn, causing it to fail to deal with large datasets. Indeed, PGM is most efficient when coupled with a marginal selection procedure that is designed to make the resulting graphical model “tree-like” (McKenna et al., 2021; Cai et al., 2021; McKenna et al., 2022). The generative network’s strong representational capacity also supports accurate feature fitting on small datasets. However, when it turns to larger and more complex datasets, its superiority in efficiency and utility will be diminished due to the greater difficulty of training high-dimensional models.
@@ 10. Conclusion and Discussion
  1. A significant trade-off exists between utility and efficiency, resulting in no single algorithm dominating the others. Machine learning-based methods, though highly efficient, generally exhibit inferior performance compared to traditional statistical methods. Nonetheless, some statistical methods, such as AIM and PrivMRF, are often time-consuming because of their heavy computation requirement, which is a trade-off for superior utility.
@@
  3. Improvement on data synthesis module remains a promising direction. We find that some feature selection methods outperform others, but existing synthesis modules exhibit significant drawbacks in different ways. For example, PGM, even though demonstrates outstanding utility, cannot work on densely selected marginals (due to high memory requests and long execution time), while generative networks face significant challenges when dealing with high-dimensional data. There remains substantial room for improvement in developing algorithms that can efficiently, reliably, and accurately generate data.
Our work not only sheds light on the strengths and limitations of current methods but also identifies several promising directions for future research. These include tailoring preprocessing strategies to algorithmic needs, developing more effective feature selection techniques, and improving the utility of machine learning models or the efficiency of some statistical methods under differential privacy. Addressing these issues and challenges can help advance this field and better support differential private data analysis.
@@ A.2. DP Mechanism
Gaussian Mechanism. Gaussian Mechanism (GM) (Mironov, 2017), which adds noise sampled from Gaussian distribution, has widely been used to achieve $(\alpha,\varepsilon)$-RDP. Specifically, Let $f$ be a vector-valued function of the input data. The Gaussian Mechanism adds i.i.d. Gaussian noise with scale $\sigma\Delta_{f}$ to each entry of $f$.
@@
Exponential Mechanism. Let $q_{r}$ be a score function for all $r\in\mathcal{R}$. Then the exponential mechanism (EM) (McKenna et al., 2022; Liu et al., 2021) outputs a candidate $r$ according to the following distribution:
@@
Gumbel Noise. Reporting noise max with Gumbel noise is a derivative mechanism from the exponential mechanism used for marginal query selection by some previous work (Vietri et al., 2020, 2022). Specifically, it outputs $i^{*}=\arg\max_{i\in[m]}\{|q_{i}(D)-a_{i}|+Z_{i}\}$, where $q_{i}$ is the marginal query function defined in  (Vietri et al., 2020, 2022); $a_{i}$ is the query answer and $Z_{i}\sim\text{Gumbel}(1/n\sqrt{2\rho})$. This mechanism has been proven to satisfy $\left(\alpha,\alpha\rho\right)$-RDP for $\forall\alpha>1$.
DP-SGD. Differential private stochastic gradient descent (DP-SGD) is the most popular way to train the model to satisfy DP. In the training process, we assume that $\mathcal{L}$ is the loss function, and we have a clipping function defined by $\text{Clip}_{C}(g)=\min\left\{1,\frac{C}{\lVert g\rVert_{2}}\right\}g$ and a Gaussian noise level $\sigma$ (Mironov et al., 2019). The DP-SGD can be expressed as
@@ B.1. PrivTree Binning
By the privacy proof of PrivTree in Zhang et al.’s work (Zhang et al., 2016), we have that if
@@
Then referring to Bun et al.’s work (Bun and Steinke, 2016), we have that if a mechanism satisfies $\varepsilon$-DP, it also satisfies $(\alpha,\frac{1}{2}\varepsilon^{2}\alpha$-RDP. Therefore we obtain that PrivTree is $(\alpha,\frac{\rho_{1}}{|V_{n}|}\alpha)$-RDP. Because we need to conduct PrivTree binning for $|V_{n}|$ rounds, by composition theorem of RDP, we have that AlgorithmAlgorithm 2 satisfies $(\alpha,\rho_{1}\alpha)$-RDP.
@@ Appendix C Missing Proof
We already have “log-sum” inequality (Csiszár, 1967), which can be expressed as
@@
where $I$ refers to mutual information (Shannon, 1948). For the right-hand side of Equation (1), applying  Lemma 1, we have the following property:
@@
$\leq$ $\sum_{A_{1},\cdots,A_{k}}\Pr[A_{1},\cdots,A_{k}]\cdot\mathbb{D}_{\text{KL}}\left.\left(\Pr[A_{i},A_{j}|A_{1},\cdots,A_{k}]\right.\;\right\|$
@@
where $I(\cdot\;|\;\cdot)$ is the conditional mutual information. Then by property of mutual information (Thomas and Joy, 2006), we have
@@ Appendix D Missing Experimental Settings
D.1. Dataset information
ACSincome and ACSemploy (Ding et al., 2021) are both drawn from 2018 national census data. The Bank dataset is on the UCI open dataset website (Moro et al., 2014), which is related to direct marketing campaigns of a Portuguese banking institution. Higgs-small (hig, 2016) dataset was produced using Monte Carlo simulations, which include different features of particles in the accelerator. Loan dataset (loa, 2020) was derived from data LendingClub issued through 2007-2014.
We make a summary of these datasets’ basic statistics in Table 5. In this table, we represent the number of records, number of attributes/numerical attributes/categorical attributes, and attributes’ domain size range of these datasets. All of these datasets have been processed to ensure that there are no missing values.
Table 5. Summary of investigated datasets.
Name #Records #Attr #Num #Cat Min/Max Domain
ACSincome (INC) (Ding et al., 2021) 55320 10 2 8 2$\sim$93
ACSemploy (EMP) (Ding et al., 2021) 37881 17 1 16 2$\sim$92
Bank (BK) (Moro et al., 2014) 45211 16 6 10 2$\sim$6024
Higgs-small (HIG) (hig, 2016) 98049 28 28 1 2$\sim$73715
Loan (LN) (loa, 2020) 134658 42 25 17 2$\sim$93995
D.2. Hyperparameters for Preprocessing algorithms
@@
D.3. Hyperparameters for Full Algorithms
@@
Upsample number $51e+$ 5 $51e+$ 5 $51e+$ 5 $51e+$ 5 $51e+$ 5
Genetic iteration $61e+$ 6 $61e+$ 6 $61e+$ 6 $61e+$ 6 $61e+$ 6
@@
Graph construction parameter $\theta$ $6$ $6$ $6$ $6$ $6$
@@
Size penalty $81e-$ 8 $81e-$ 8 $81e-$ 8 $81e-$ 8 $81e-$ 8
Max marginal attributes number $6$ $6$ $6$ $6$ $6$
Max clique size $71e+$ 7 $71e+$ 7 $71e+$ 7 $71e+$ 7 $71e+$ 7
@@
Random projection number $62e+$ 6 $62e+$ 6 $62e+$ 6 $62e+$ 6 $62e+$ 6
Categorical optimization rate $33e-$ 3 $33e-$ 3 $33e-$ 3 $33e-$ 3 $33e-$ 3
Numerical optimization rate $36e-$ 3 $36e-$ 3 $36e-$ 3 $36e-$ 3 $36e-$ 3
Top q $5$ $5$ $5$ $5$ $5$
Categorical optimization step $1$ $1$ $1$ $1$ $1$
Numerical optimization step $3$ $3$ $3$ $3$ $3$
Upsample rate $10$ $10$ $20$ $20$ $40$
@@
Random feature dimension $32e+$ 3 $32e+$ 3 $32e+$ 3 $32e+$ 3 $32e+$ 3
Mini batch rate $25e-$ 2 $25e-$ 2 $25e-$ 2 $25e-$ 2 $25e-$ 2
Epoch number $31e+$ 3 $31e+$ 3 $31e+$ 3 $31e+$ 3 $31e+$ 3
Learning rate $21e-$ 2 $21e-$ 2 $21e-$ 2 $21e-$ 2 $21e-$ 2
@@
Denoiser layer number $2$ $2$ $2$ $2$ $2$
@@
D.4. Hyperparameters for Different Feature Selection Algorithms
Table 15. RAP++ Selection Hyperparameters
Hyperparameter INC EMP BK HIG LN
Top q $3$ $3$ $3$ $3$ $3$
Selection step $4$ $6$ $6$ $7$ $10$
Selection budget rate $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
Marginal budget rate $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
Table 16. PrivSyn Selection Hyperparameters
Hyperparameter INC EMP BK HIG LN
Selection budget rate $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
1-way marginal budget rate $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
2-way marginal budget rate $0.8$ $0.8$ $0.8$ $0.8$ $0.8$
D.5. Hyperparameters for Different Synthesis Algorithms
Table 17. Relaxed Projection Hyperparameters
Hyperparameter INC EMP BK HIG LN
Random projection number $62e+$ 6 $62e+$ 6 $62e+$ 6 $62e+$ 6 $62e+$ 6
Optimization rate $35e-$ 3 $35e-$ 3 $35e-$ 3 $35e-$ 3 $35e-$ 3
Optimization step $100$ $170$ $160$ $280$ $420$
Table 18. Generative Network Hyperparameters
Hyperparameter INC EMP BK HIG LN
Learning rate $31e-$ 3 $31e-$ 3 $31e-$ 3 $31e-$ 3 $31e-$ 3
Synthesis size $1024$ $1024$ $1024$ $1024$ $1024$
Max training iteration $50$ $50$ $100$ $100$ $1500$

### v2

@@ 9.3. Module Comparison
Comparison of Synthesis Modules. We compare the synthesis algorithms of all six algorithms: GUM, PGM, relaxed projection (RP), genetic algorithm (GA), and generative network (GN). PGM is employed by both AIM and PrivMRF. To avoid multiple fitting steps from adaptive selection and control the running time, we use the PrivSyn selection algorithm as the common marginal selector. The results are shown in Figure 8. GUM and PGM are conducted on CPU, and the remaining works are on GPU.
Among these methods, PGM and the generative network achieve the best overall accuracy and stability in fitting features on small datasets. PGM benefits from its expressive structure, which, as indicated by Equation 2, infers marginals without refitting existing ones, thus minimizing compounding errors. However, the PGM’s fitting process is quite slow, due to the densely selected marginals by PrivSyn, causing it to fail to deal with large datasets. Indeed, PGM is most efficient when coupled with a marginal selection procedure that is designed to make the resulting graphical model “tree-like” (mckenna2021winning; cai2021data; mckenna2022aim). The generative network’s strong representational capacity also supports accurate feature fitting on small datasets. However, when it turns to larger and more complex datasets, its superiority in efficiency and utility will be diminished due to the greater difficulty of training high-dimensional models.
@@ 10. Conclusion and Discussion
  1. A significant trade-off exists between utility and efficiency, resulting in no single algorithm dominating the others. deep learning-based methods, though highly efficient, generally exhibit inferior performance compared to traditional statistical methods. Nonetheless, under current implementations, some statistical methods, such as AIM and PrivMRF, are often time-consuming, which is a trade-off for superior utility.
@@
  3. Improvement on data synthesis module remains a promising direction. We find that some feature selection methods outperform others, but existing synthesis modules exhibit significant drawbacks in different ways. For example, even though PGM demonstrates outstanding utility, it cannot work on densely selected marginals (due to high memory requests and long execution time), while generative networks face significant challenges when dealing with high-dimensional data. There remains substantial room for improvement in developing algorithms that can efficiently, reliably, and accurately generate data.
Table 9. Estimated cost (US dollar) of running different algorithms.
Method PrivSyn PrivMRF AIM RAP++ Private-GSD GEM DP-MERF TabDDPM
Total Cost \$ 0.2 \$ 1.1 \$ 5.7 \$ 68.1 \$ 17.0 \$ 0.7 \$ 0.1 \$ 2.1
Our work not only sheds light on the strengths and limitations of current methods but also identifies several promising directions for future research. These include tailoring preprocessing strategies to algorithmic needs, developing more effective feature selection techniques, and improving the utility of deep learning models or the efficiency of some statistical methods by more efficient implementation or GPU acceleration. In addition to these research opportunities, there are still certain other aspects, though not investigated in this work, worth exploring. For example, beyond the time efficiency perspective, dollar cost is a more realistic factor in algorithm selection. Based on our experimental setups and hardware, we roughly set the cost at \$0.8/hr for GPU usage, and \$0.2/hr for CPU usage (the actual prices may vary due to many factors such as dataset complexity, which can affect our selection of hardware, and price fluctuations of cloud service (price1; price2; price3; price4), etc.). We summarize the total price consumption of all algorithms during the overall comparison in Table 9. We can observe that those methods with lower time efficiency tend to incur higher costs.
Finally, we discuss the scalability of our framework. It offers a broader perspective for analyzing existing synthesis algorithms, but we also need to recognize its limitations. For example, we regard gradients as features to explain those methods based on DP-SGD. But the gradient feature itself is not interpretable enough. Besides, some “model-level DP” methods may fall outside our three-step workflow. For instance, PATE-GAN (jordon2018pate) uses multiple teacher discriminators to achieve DP and trains the student discriminator on them to further guide the data generator. In this case, our framework may lack corresponding explanatory power and require further extension to explain these methods.
Acknowledgements.
We thank anonymous reviewers for their valuable comments. This paper was partially supported by NSF CNS-2220433, NSF OAC-2319988, and a CCI award.
@@ A.2. DP Mechanism
Gaussian Mechanism. Gaussian Mechanism (GM) (mironov2017renyi), which adds noise sampled from Gaussian distribution, has widely been used to achieve $(\alpha,\varepsilon)$-RDP. Specifically, Let $f$ be a vector-valued function of the input data. The Gaussian Mechanism adds i.i.d. Gaussian noise with scale $\sigma\Delta_{f}$ to each entry of $f$.
@@
Exponential Mechanism. Let $q_{r}$ be a score function for all $r\in\mathcal{R}$. Then the exponential mechanism (EM) (mckenna2022aim; liu2021iterative) outputs a candidate $r$ according to the following distribution:
@@
Gumbel Noise. Reporting noise max with Gumbel noise is a derivative mechanism from the exponential mechanism used for marginal query selection by some previous work (vietri2020neworacleefficientalgorithmsprivate; vietri2022private). Specifically, it outputs $i^{*}=\arg\max_{i\in[m]}\{|q_{i}(D)-a_{i}|+Z_{i}\}$, where $q_{i}$ is the marginal query function defined in  (vietri2020neworacleefficientalgorithmsprivate; vietri2022private); $a_{i}$ is the query answer and $Z_{i}\sim\text{Gumbel}(1/n\sqrt{2\rho})$. This mechanism has been proven to satisfy $\left(\alpha,\alpha\rho\right)$-RDP for $\forall\alpha>1$.
DP-SGD. Differentially private stochastic gradient descent (DP-SGD) is the most popular way to train the model to satisfy DP. In the training process, we assume that $\mathcal{L}$ is the loss function, and we have a clipping function defined by $\text{Clip}_{C}(g)=\min\left\{1,\frac{C}{\lVert g\rVert_{2}}\right\}g$ and a Gaussian noise level $\sigma$ (mironov2019r). The DP-SGD can be expressed as
@@ B.1. PrivTree Binning
By the privacy proof of PrivTree in Zhang et al.’s work (zhang2016privtree), we have that if
@@
Then referring to Bun et al.’s work (bun2016concentrated), we have that if a mechanism satisfies $\varepsilon$-DP, it also satisfies $(\alpha,\frac{1}{2}\varepsilon^{2}\alpha$-RDP. Therefore we obtain that PrivTree is $(\alpha,\frac{\rho_{1}}{|V_{n}|}\alpha)$-RDP. Because we need to conduct PrivTree binning for $|V_{n}|$ rounds, by composition theorem of RDP, we have that AlgorithmAlgorithm 2 satisfies $(\alpha,\rho_{1}\alpha)$-RDP.
@@ Appendix C Missing Proof
We already have “log-sum” inequality (csiszar1967information), which can be expressed as
@@
where $I$ refers to mutual information (shannon1948mathematical). For the right-hand side of Equation (1), applying  Lemma 1, we have the following property:
@@
[…]
@@
where $I(\cdot\;|\;\cdot)$ is the conditional mutual information. Then by property of mutual information (thomas2006elements), we have
@@ Appendix D Missing Experimental Settings
D.1. Supplementary Dataset information
ACSincome and ACSemploy (ding2021retiring) are both drawn from 2018 national census data. The Bank dataset is on the UCI open dataset website (bank_marketing_222), which is related to direct marketing campaigns of a Portuguese banking institution. Higgs-small (higgsdata) dataset was produced using Monte Carlo simulations, which include different features of particles in the accelerator. Loan dataset (loandata) was derived from data LendingClub issued through 2007-2014.
D.2. Implementations
D.3. Hyperparameters for Preprocessing algorithms
@@
D.4. Hyperparameters for Full Algorithms
@@
Upsample number $1e+5$ $1e+5$ $1e+5$ $1e+5$ $1e+5$
Genetic iteration $1e+6$ $1e+6$ $1e+6$ $1e+6$ $1e+6$
@@
Graph construction parameter $\theta$ 66 66 66 66 66
@@
Size penalty 1e81e-8 1e81e-8 1e81e-8 1e81e-8 1e81e-8
Max marginal attributes number 66 66 66 66 66
Max clique size $1e+7$ $1e+7$ $1e+7$ $1e+7$ $1e+7$
@@
Random projection number $2e+6$ $2e+6$ $2e+6$ $2e+6$ $2e+6$
Categorical optimization rate 3e33e-3 3e33e-3 3e33e-3 3e33e-3 3e33e-3
Numerical optimization rate 6e36e-3 6e36e-3 6e36e-3 6e36e-3 6e36e-3
Top q 55 55 55 55 55
Categorical optimization step 11 11 11 11 11
Numerical optimization step 33 33 33 33 33
Upsample rate $10$ $10$ 2020 2020 $40$
@@
Random feature dimension $2e+3$ $2e+3$ $2e+3$ $2e+3$ $2e+3$
Mini batch rate 5e25e-2 5e25e-2 5e25e-2 5e25e-2 5e25e-2
Epoch number $1e+3$ $1e+3$ $1e+3$ $1e+3$ $1e+3$
Learning rate 1e21e-2 1e21e-2 1e21e-2 1e21e-2 1e21e-2
@@
Denoiser layer number 22 22 22 22 22
@@
D.5. Hyperparameters for Different Feature Selection Algorithms
Table 18. RAP++ Selection Hyperparameters
Hyperparameter INC EMP BK HIG LN
Top q 33 33 33 33 33
Selection step 44 66 66 77 $10$
Selection budget rate $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
Marginal budget rate $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
Table 19. PrivSyn Selection Hyperparameters
Hyperparameter INC EMP BK HIG LN
Selection budget rate $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
1-way marginal budget rate $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
2-way marginal budget rate $0.8$ $0.8$ $0.8$ $0.8$ $0.8$
D.6. Hyperparameters for Different Synthesis Algorithms
Table 20. Relaxed Projection Hyperparameters
Hyperparameter INC EMP BK HIG LN
Random projection number $2e+6$ $2e+6$ $2e+6$ $2e+6$ $2e+6$
Optimization rate 5e35e-3 5e35e-3 5e35e-3 5e35e-3 5e35e-3
Optimization step $100$ $170$ $160$ $280$ $420$
Table 21. Generative Network Hyperparameters
Hyperparameter INC EMP BK HIG LN
Learning rate 1e31e-3 1e31e-3 1e31e-3 1e31e-3 1e31e-3
Synthesis size $1024$ $1024$ $1024$ $1024$ $1024$
Max training iteration $50$ $50$ $100$ $100$ $1500$
@@ D.6. Hyperparameters for Different Synthesis Algorithms
Conditional Query Error $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$
PrivSyn 0.063 0.019 0.017 0.048 0.035 0.030 0.034 0.033 0.030 0.034 0.029 0.024 0.070 0.058 0.052
PrivMRF 0.014 0.012 0.009 0.022 0.013 0.010 0.031 0.026 0.024 0.025 0.026 0.025 0.054 0.053 0.051
RAP++ 0.029 0.021 0.020 0.039 0.026 0.019 0.072 0.064 0.063 0.099 0.098 0.104 0.108 0.098 0.092
AIM 0.030 0.013 0.010 0.028 0.011 0.008 0.033 0.022 0.022 0.028 0.025 0.025 0.058 0.057 0.050
Private-GSD 0.024 0.016 0.011 0.027 0.014 0.009 0.045 0.039 0.037 0.110 0.110 0.107 0.091 0.093 0.096
GEM 0.025 0.026 0.026 0.042 0.028 0.026 0.055 0.053 0.051 0.068 0.069 0.073 0.077 0.077 0.077
DP-MERF 0.308 0.308 0.311 0.316 0.314 0.313 0.210 0.215 0.238 0.101 0.099 0.104 0.416 0.455 0.426
TabDDPM 0.215 0.092 0.083 0.268 0.072 0.079 0.131 0.136 0.144 0.566 0.109 0.109 0.143 0.143 0.140
Appendix E Supplementary Experiment Results
E.1. More Results of Algorithm Comparison
We have provided some important metric results of algorithms’ utility in Section 9.1. Here we list some other detailed evaluations, such as AUC, accuracy for machine learning efficiency, and running time for time efficiency. Notably, these results do not necessarily influence our conclusion. Thus, we present them in the appendix.
E.2. Detailed Results of Preprocessing Influence
Table 23. Comparison of fidelity error and Cramer’s V measure error on Bank dataset under $\varepsilon=0.2$.
Method Fid Err. Cramer Err. Method Fid Err. Cramer Err.
Privsyn 0.24 0.10 Private-GSD 0.52 0.11
PrivMRF 0.13 0.07 GEM 0.24 0.10
RAP++ 0.43 0.10 DP-MERF 0.47 0.11
AIM 0.11 0.06 TabDDPM 0.77 0.13